OpenStudy (sydthekid913):

RAINBOW KITTY UNIORN AND MEDAL TO THE BRILLIANT MIND WHO CAN ANSWER THIS ^.^ Determine which polynomial has (x + 4) as a factor. A. 2003-16-03-00-00_files/i0210000.jpg B. 2003-16-03-00-00_files/i0210001.jpg C. 2003-16-03-00-00_files/i0210002.jpg D. 2003-16-03-00-00_files/i0210003.jpg

OpenStudy (sydthekid913):

whoops

OpenStudy (sydthekid913):

A. x^2 - 11x + 28 B. x^2 - 6x + 8 C. x^2 + 7x + 12 D. x^2 - 4x + 16

OpenStudy (luffingsails):

(x+4)(x+3) = x^2 + 7x + 12. So, C.

OpenStudy (sydthekid913):

why couldn't it be B tho??

OpenStudy (luffingsails):

What is your other term? Meaning (x + ?)(x + 4) = x^2 - 6x + 8. There's no real answer to fill in the ? with.

OpenStudy (sydthekid913):

couldn't the other term be 2? the info above is all they gave me :\

OpenStudy (luffingsails):

If I put 2 in for the question mark, I get: (x+2)(x+4) = x^2 + 6x + 8.... but it has to be -6x not +6x. See?

OpenStudy (sydthekid913):

Oh thank you so much!!!

OpenStudy (luffingsails):

You're welcome.